Write the expression as a complex number in standard form.
a) 8/1+i
b) -5-3i/4i
c) 2+5i/5+2i

Answer by venugopalramana(3286)

Write the expression as a complex number in standard form.
a) 8/1+i = [8*(1-I)]/[(1+I)(1-I)]=[8-8I]/(1+1)=4-4I
b) -5-3i/4i =-[5+3I][I]/[5I*I]=[-5I+3]/(-5)=(-3/5)+I
c) 2+5i/5+2i = [(2+5I)(5-2I)]/[(5+2I)(5-2I)]=[10-4I+25I+10]/[25+4]=[20-21I]/29
=(20/29)-(21/29)I